# T2Vid: Translating Long Text into Multi-Image is the Catalyst for Video-LLMs

### v1

@@ 5.2.3 A Look at Instruction Diversity
Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT. 5,000 instructions are sampled from each dataset. ShareGemini exhibits some clear clusters, while Video-ChatGPT is relatively more diverse.
@@
Prior studies have underscored the importance of instruction diversity for fine-tuning LLMs [46] and image-LLMs [42]. These findings prompt us to conduct an inspection of data in this aspect. Thus, we follow previous approaches [36, 45] to visualize the diversity of instructions in the training corpus. Specifically, we randomly sample 5,000 instructions from ShareGemini and Video-ChatGPT, respectively. Then, the instructions are embedded and visualized using the t-SNE technique, as shown in Fig. 4. In the figure, the distribution of ShareGemini exhibits 9 clear clusters. This is because the dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail.” In contrast, the distribution of Video-ChatGPT is more diverse, as it includes not only video summarization but also questions related to specific video content. Overall, the instructions for these video samples are not diverse enough.
6 Our Methods
Since currently available video data are limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity using synthetic data.
@@ 6 Our Methods
A rich source of instruction data lies in the text domain, and they can effectively enrich the diversity of instruction. Nevertheless, there is inherently a modality gap between the text and visual domain. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates our overall idea. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments and embed the text of each segment into a sequence of images. For example, the context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g. an inquiry to summarize the given paragraphs and the corresponding synopsis. After the transformation, the structure of these instruction data is exactly the same as the video counterparts, and we can incorporate the synthetic data into the video training corpus without extra processing.
Figure 6: t-SNE plot of instruction distribution after applying our proposed methods. By incorporating text data, i.e. LongAlpaca [5] and LongQLoRA [5], the overall distribution becomes more diverse. 5,000 instructions are sampled from each dataset.
@@
On a related note, LLaMA-VID [20] and LongVA [43] also utilize text data to assist video understanding, and we underscore the differences here: (1) Different objectives: LLaMA-VID and LongVA use text data in the fine-tuning and continued pre-training stage, respectively, in order to expand the context window for long video understanding. In contrast, our methods mainly utilize the diversity of instructions in the text data to facilitate efficient fine-tuning. (2) Different processing methods: LLaMA-VID and LongVA directly use the pure text data in training, while our methods transform text samples into video-like samples and unify the training format.
@@
Table 3: Accuracy comparisons of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K). The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
Video-MME MVBench TempCompass Avg.
Short Medium Long Overall
Text-aided Models
LLaMA-VID [20] 40.6 40.2 34.9 38.6 41.9 45.6 42.0
LongVA [43] 61.1 48.8 45.4 51.8 50.7 56.9 53.1
MiniCPM-8B [37]
Zero-shot 58.3 46.1 40.1 48.2 42.9 49.1 46.7
200K data 60.3 47.3 44.7 50.8 48.0 54.7 51.2
30K data 61.9 51.7 45.6 53.0 48.4 56.8 52.7
Idefics3-8B [17]
Zero-shot 61.3 47.0 45.1 51.2 49.6 55.9 52.2
200K data 62.0 53.1 44.9 53.3 50.7 62.9 55.6
30K data 67.1 53.8 48.0 56.3 51.6 62.3 56.7
Table 4: The performance of different models/settings on three video benchmarks. Our methods use only 15% of the total sample size compared to the full video sets (200K) for fine-tuning and achieve similar or even superior performance. The highest performance of a model among all settings is highlighted in bold. LLaMA-VID and LongVA also use pure text data but without transformation, which is different from our methods and objectives.
7 Evaluation on Proposed Methods
7.1 Ablation on Different Data Compositions
@@
The effectiveness of our proposed methods. As shown in Tab. 3, comparing the first three rows, we can find that when using the same amount of video samples, training only with ShareGemini is not as effective as using more diverse data compositions. Meanwhile, using the same amount of data, our proposed scheme (the 4th row) achieves the best performance. Moreover, when compared with the full 200K data fine-tuning setting, our proposed scheme uses much fewer data samples (only 15%) to achieve comparable performance, and the training cost reduces from 276.8 GPU hours to 33.6 GPU hours, making an 8.2x speedup. The overall results demonstrate the importance of instruction diversity and the effectiveness of our proposed methods.
@@
Notably, replacing the synthetic data with the original pure text counterpart achieves an overall inferior performance. We hypothesize that this is due to the inherent domain gap between vision and text. Thus, to simulate the structure of video sequences, transcribing long text into images might be necessary.
@@
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab. 3). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results to further illustrate this point.
7.2 Generalization Ability of Our Proposed Scheme
We further verify the generalization ability of the proposed scheme and evaluate our methods on more image-LLMs of larger parameter sizes, including MiniCPM-8B and Idefics3-8B, across different benchmarks, as shown in Tab. 4.
General effectiveness of the proposed scheme. The results in the table suggest that on both models, though fine-tuning with fewer samples, our proposed scheme achieves performance on par with or better than fine-tuning with the complete video datasets. We also compare with related methods, LLaMA-VID and LongVA, which also utilize text data but in the original form of pure text, in order to expand the context window. In contrast, our methods excel in enriching the instructions via synthesizing video-like data from text data.
Boost in long video understanding. On the long video sets of the Video-MME benchmark, our methods provide consistent gains compared to the full data fine-tuning setting, though we use video samples (most no longer than 3.5 minutes in the training sets) of lengths much shorter than the test sets (30-60 minutes). The results indicate that our proposed methods can assist models in understanding long videos.
@@ 8 Conclusion
This paper revisits the two mainstream approaches of utilizing image-LLMs to perform video understanding. A chain of investigations reveals possible drawbacks in both approaches, underscoring the bottlenecks of zero-shot inference and the easily ignored aspects of fine-tuning. Among these discoveries, the limited instruction diversity of current video data is notable, leading to downgraded learning efficiency. Aimed at this specific issue, we develop an efficient method, dubbed T2Vid, to enrich the instructions in the training corpus. Comprehensive evaluations demonstrate the effectiveness of the proposed method. We hope the findings from this paper can spark deeper thinking about using MLLMs for video understanding and how to curate video data of higher quality.
Appendix A Implementation Details
@@ A.1 Answer Judging
Appendix B Additional Results in Experiments
B.1 Results on Different Frame Counts
Figure 7: Accuracy of MiniCPM-8B and Idefics3-8B on the Video-MME benchmark when inference with different numbers of video frames. As input frames increase, the performance of both models stays relatively stable and does not improve much. “max training frames” denotes the max number of patches used for a single image when training; “max training context” denotes the maximum context window set when training; “max LLM context” denotes the context window of the original LLM backbone before further multimodal training.
We present the performance of different input frame numbers of MiniCPM-8B [37] and Idefics3-8B [17] in Fig. 7. As input frames increase, the performance within the context window of the LLM backbone is more stable than the InterVL-4B, which is of a smaller parameter size. Nevertheless, the performance still does not improve much as more frames are available, again indicating that the effective context is limited. Thus, simply expanding the LLM context window for image-LLMs to facilitate long video understanding would not be very practical. As in the main paper, we attribute this phenomenon to the lack of capabilities of image-LLMs to truly understand the temporal relationships among frames. Additional case studies related to this topic are available in the following section.
B.2 Case Study
In the main paper, we identify the drawback of image-LLMs generally failing to understand the temporal relationships among video frames. In this section, we supplement more cases and analyses on this topic.
Figure 8: A case study using the prompt “Describe this video.”, where the video is a documentary of the Arctic environment. Compared with the fine-tuned model, the zero-shot case reasons with and describes video frames without understanding the temporal correlations. This pattern follows the common behavior of the image model. The keyframes are highlighted with colors (except for red, which is used to indicate hallucinations). The numbers on the upper-left corner of each frame indicate frame order. Zoom in for better visibility.
In case 1 (Fig. 8), we list a case where we instruct models to describe the video, while the image-LLM fails to understand the separate scenes and temporal relationships but instead focuses on several specific frames and describes rather independently. In contrast, the model fine-tuned with video data can more accurately distinguish the actions happening in the video and understand the connections in a temporally coherent way.
Figure 9: A case study designed to test the temporal reasoning capabilities of models. In the video, three polar bears show up at the beginning, and in the end, the scene changes into a video shot of two polar bears. The image-LLM fails to recognize the change through zero-shot prompting. Note that we have verified that the image-LLM can correctly recognize the number by inputting a single frame separately.
In case 2 (Fig. 9), we ask the two models to identify the temporal changes in object counts. The fine-tuned model correctly recognizes the changes in the number of polar bears as scenes change in the video clip, while the image-LLM fails to do so and seems to give the answer according to the last few frames. This is expected since image-LLMs are trained with static images rather than dynamic videos, where the latter inject temporal concepts into models.
@@ B.2 Case Study
If the meaning of all options are significantly different from the answer, output Z. Your should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else. Here are two examples.
@@
Question: What is the main object in image?
@@
Question: What is the main object in image?
@@
Now here is the question, options, and the answer, you should match and give me the option letter:

### v3

@@ 3.3.1 Low Learning Efficiency Issue
Figure 4: t-SNE plot of instruction distribution of video datasets - ShareGemini and Video-ChatGPT. 5,000 instructions are sampled from each dataset. The instruction diversity of these two datasets is relatively limited, leading to low learning efficiency.
Our experiments start with scaling up the training data volume and evaluating the video understanding performance on different general video understanding benchmarks. The results are shown in Fig. 3. In general, training either with video caption data (ShareGemini), instruction data (Video-ChatGPT), or a mix of both can boost the image-LLM’s video understanding performance. Meanwhile, increasing the training volume brings additional gains in accordance with the data scaling law. However, the gains from scaling up quickly reach a plateau. For instance, on the Video-MME benchmark, when training with mixed data, 30K samples improve overall accuracy by 3.1 points, while 100K samples only add another 0.5 points, resembling a logarithmic growth. In view of this quick and early saturation, the learning efficiency with these video datasets can be quite limited. The phenomenon also suggests that there could be high redundancy in the training corpus, and it is possible that we may use less data to achieve a performance comparable to or even better than training with more data samples.
3.3.2 Probing of Instruction Diversity
@@
Previous results prompt us to explore the reason for such low learning efficiency. Inspired by prior studies, which have underscored the importance of instruction diversity for fine-tuning LLMs [37] and image-LLMs [38], we conduct an inspection of training data in this aspect. Specifically, we follow previous approaches [39, 40] to visualize the distribution of instructions in the training corpus. 5,000 instructions are sampled from ShareGemini and Video-ChatGPT, respectively. Then, the instructions are embedded and visualized using the t-SNE technique, as shown in Fig. 4. Overall, the instruction distribution of these two datasets is not diverse enough, which leads to a low data efficiency: The distribution of ShareGemini exhibits 9 clear clusters in the figure, indicating very similar instructions. This is because this dataset samples from a fixed pool of 9 templates as instructions, each of which is a variant of “Describe this video in detail”. On the other hand, the distribution of Video-ChatGPT seems relatively more diverse, as it includes specific questions related to video content and details besides video summarization. Nevertheless, the instruction diversity is still low due to the nature of self-instruction and a few fixed task-specific prompting templates for data curation.
4 Methods
4.1 Design Concept
@@
Since currently available video data can be limited in instruction diversity, and annotating high-quality video data is costly, we aim to expand the instruction diversity by incorporating new synthetic data. A rich source of instruction data lies in the text domain, and they can effectively complement the vision domain. Nevertheless, there is inherently a modality gap between the text and visual domains. To better utilize these data, we bridge the modality gap by synthesizing images with the text. Fig. 5 illustrates our overall workflow for synthesis of data.
Figure 6: t-SNE plot of instruction distribution after applying our proposed methods. By incorporating text data, i.e. LongAlpaca [41] and LongQLoRA [41], the overall distribution becomes more diverse.
@@
4.2 Implementation Details
Each text sample is a (long-context, question, answer) triplet. For example, the long context can be a section of a book or an academic paper, while the instruction and the answer are centered around the context, e.g. an inquiry to give a synopsis, or questions related to the paper. After the data transformation process, each sample is a video-like (images, question, answer) triplet, where long-context information is transformed into a series of images, and the question and answer stay unchanged.
The key to the data synthesis procedure is synthesizing images with pure text. Specifically, for each (long-context, question, answer) triplet, we divide the context information into multiple segments according to word counts (set to 115 empirically) using an open-sourced NLP toolkit [^2] https://www.nltk.org/. These text chunks are then transformed into a sequence of images separately. Specifically, each text chunk is embedded into a blank image with a white background. This is achieved using bitmap font with the ImageFont module of Pillow library [^3] https://pillow.readthedocs.io/en/stable/. Each image is 448x448 pixels in size, and the font is 20 pt large, black color, Arial Regular type. We use a bounding box to control the layout, leaving a margin of 20 pixels on each side, so each line of the text has roughly the same width. After the transformation, the structure of these synthesized data is exactly the same as that of video samples, and we can seamlessly incorporate the synthetic data into the video training corpus.
5 Evaluation on Proposed Methods
This section includes experimental results and discussions of our proposed method, including (1) a comparison with mainstream methods, (2) an ablation study on data mixes, and (3) an examination of key properties, including data scaling performance and gains in long video understanding.
Methods Size Frames Short Medium Long Overall
Proprietary Models
GPT-4V [42] N/A 10 70.5 55.8 53.5 59.9
Claude 3.5 Sonnet [43] N/A 20 71.0 57.4 51.2 60.0
GPT-4o [44] N/A 384 80.0 70.3 65.3 71.9
Gemini 1.5 Pro [45] N/A 1fps 81.7 74.3 67.4 75.0
Open-Source Models
VideoChat2 [9] 7B 16 48.3 37.0 33.2 39.5
Video-LLaVA [13] 7B 8 45.3 38.0 36.2 39.9
Chat-UniVi-v1.5 [14] 7B 64 45.7 40.3 35.8 40.6
VideoLLaMA 2 [8] 7B 16 56.0 45.4 42.1 47.9
VITA [46] 8x7B 32 65.9 52.9 48.6 55.8
Kangaroo [47] 8B 64 66.1 55.3 46.6 56.0
VITA-1.5 [48] 7B 16 67.0 54.2 47.1 56.1
FT w/ InternVL [7]
Zero-shot 3.8B 64 61.3 51.8 44.3 52.5
200K video data 3.8B 64 66.7 54.2 48.1 56.3
Sparrow (30K hybrid data) 3.8B 64 67.0 53.7 49.3 56.7
Table 1: Accuracy comparisons of different methods on the Video-MME benchmark. Performance is ranked in ascending order regarding overall performance (The models are grouped according to open-source or not). Our method uses only 15% of the total sample size compared to the full volume video datasets (200K) for fine-tuning and achieves comparable performance. 30K hybrid data comprise 20K data sampled from video datasets and 10K synthesized from our method. Bold digits indicate the best performance within each group.
5.1 Comparison with Mainstream Methods
We compare the results with some representative proprietary models, including GPT-4V [42], Claude 3.5 Sonnet [43], GPT-4o [44], Gemini 1.5 Pro [45], and open-source video-LLMs of similar LLM parameter size, including Video-LLaVA [13], VideoChat2 [9], Chat-UniVi-v1.5 [14], VideoLLaMA 2 [8], VITA [46], and Kangaroo [47], as summarized in Tab. 1.
The table results show that, through zero-shot inference, the image-LLM Intern-VL already outperforms a variety of video-LLMs with larger LLM parameter sizes. This might be due to the rich pre-trained knowledge embedded in the model parameters since the image-LLM has been trained with large-scale and high-quality image-text data. The vision prior lays a strong foundation for further video fine-tuning, where models learn temporal and causal concepts from activities, events, etc. The model fine-tuned with full video datasets achieves an overall gain of 3.8 points on the image-LLM, closing the gap between open-source models and proprietary ones.
Notably, our methods use only 15% of the total sample size compared to the full volume (200K) for fine-tuning and achieve comparable performance. This result suggests the high data efficiency of our proposed scheme since mixing in synthetic samples mitigates the low instruction diversity issue illustrated in the earlier section.
5.2 Ablation on Different Data Compositions
@@
Table 2: Results of different data compositions on the Video-MME benchmark. Our proposed scheme achieves an overall performance superior to other data mixes of the same amount (30K) and even more data (200K). The Zero-shot and 200K full data fine-tuning settings are listed for comparison.
@@
[…]
@@
Notably, replacing the synthetic data with the original pure text counterpart achieves an overall inferior performance. We hypothesize that this is due to the inherent domain gap between vision and text. Thus, to simulate the structure of video frame sequences, transcribing long text into images is necessary.
Figure 7: Scaling performance with different data volumes and types on general video benchmarks. We consider four settings, i.e. fine-tuning with video data only, our hybrid data (video data mixed with our synthetic samples with a ratio of 2:1:212:12 : 1), TOPA data, and T3 data. By default, video data (the caption dataset and the instruction dataset) are mixed with a 1:$1:11$ : 1 sampling ratio. Training samples are measured in K, where 0 sample indicates zero-shot inference. Purple-shaded area indicates gains of our methods over the video data baseline.
@@
Can synthetic data help models understand longer videos? Interestingly, in the training stage, we only utilize synthetic samples of long multimodal context instead of authentic long video samples. However, on the long video benchmark set, our proposed method still achieves a score that is 1.2 points higher than the full data training (as shown in the 4th row compared to 200K full data in Tab. 2). This result suggests that fine-tuning with a long multimodal context can enhance the comprehension of longer videos. In the following section, we will present additional results and discussions to illustrate this point further.
[…]
Baseline 0 24 40.1 40.0 44.5 41.6
30 24 44.7 39.7 45.4 43.3
60 24 46.2 42.7 46.2 45.1
100 24 46.7 44.1 45.3 45.3
Our Method 30 24 45.6(+0.9) 48.7(+9.0) 51.4(+6.0) 48.5(+5.2)
60 24 46.2 51.2(+8.5) 53.2(+7.0) 50.2(+5.1)
100 24 48.7(+2.0) 50.1(+6.0) 57.0(+11.7) 51.9(+6.6)
Table 3: Long video understanding performance with different data volumes and types. The Baseline method adopts only video data, while Our Method utilizes a mix of video data and synthetic samples attained from our method. In the Baseline group, 0 sample indicates zero-shot inference. The performance gains are calculated relative to the video data fine-tuning baseline trained with the same number of samples.
5.3 Examination of Key Properties
In this section, we further examine key properties of our proposed method, including general effectiveness, scaling performance, and effectiveness when adopted in long video understanding scenarios.
5.3.1 General Effectiveness and Scaling Performance
We further verify the proposed scheme’s effectiveness by evaluating our methods on another image-LLM of larger parameter size, i.e. MiniCPM-8B, across different benchmarks. Specifically, through scaling up with different volumes and types of data, we compare our methods against a pure video data baseline, as well as other relevant methods, including TOPA and T3. Both TOPA and T3 first translate vision information into text, such as captions and relations between objects. Then, synthetic text QAs are constructed to simulate video reasoning samples, aiming to transfer temporal reasoning capabilities from text to video. Note that since the original data format of TOPA is largely different from the current paradigm, we design a template to adapt them to the instruction data format (More details are available in Appendix A.2). The results are summarized in Fig. 7.
General effectiveness. When using the same amount of training samples, our methods almost always outperform other methods by a clear margin in all the evaluated benchmarks. Specifically, when using 30K samples, our method achieves an overall accuracy of 52.7, surpassing the baseline by 3.9 points. Notably, it even outperforms the baseline trained with 100K samples by 1.7 points. Similarly, on the MVBench benchmark, with 100K samples, our method attains a 4.3-point absolute gain over the baseline. Overall, the superior performance on different benchmarks showcases the general effectiveness of our proposed method.
Scaling performance. A clear issue with other methods is that they are more prone to performance saturation when scaling up the data budget. For instance, the baseline method uses 60K samples to improve the Video-MME benchmark by 3.1 points, while using 100K samples only achieves another 0.9 points of absolute gains. In contrast, our proposed scheme shows more stable and consistent improvements when scaling up the data volumes compared with other methods. This suggests that ensuring a more diverse distribution of instruction in video training is very important, or otherwise, the learning efficiency may downgrade. To avoid such circumstances, the proposed data augmentation method can be an effective approach.
Discussion: Can we scale up only with synthetic textual samples? An intriguing and highly relevant question is whether we can scale up the synthetic samples without using any real video samples. Since the text is more compact and less redundant than a whole video, training in this way is more economical. Unfortunately, the empirical results show that this is probably unfeasible. As shown in Fig. 7, scaling with synthetic text data (TOPA and T3) shows undesirable characteristics, i.e. this approach can easily reach the saturation point or even slightly downgrade. Other critical issues include the modality gap and special processing of videos in various domains (such as egocentric videos and movies). Besides, since text suffers inevitably from information loss when translated from videos, text data might be better used as a supplement to videos, which helps inject temporal reasoning language prior into the LLM backbones (similar to TOPA or T3) or mixing with video data as a regularization method (as our method does).
Figure 8: Long video understanding performance evaluated with a different number of frames.Video context” denotes the max number of video frames used in training, i.e. 24 frames; “LLM context” denotes the context window of the LLM backbone. Purple-shaded area indicates performance gains of our proposed method over the video data baseline.
5.3.2 Long Video Understanding Performance
We adopt tailored benchmarks to evaluate long video understanding capabilities, including LongVideoBench [49], MLVU-M [50], and the long video set of Video-MME, and report the performance on evaluation sets for the former two benchmarks. Our study focuses on two aspects: (1) performance improvement in terms of long video understanding compared with the video fine-tuning baseline and (2) frame number (multimodal context) generalization ability in the inference stage.
Performance improvement w.r.t. training sample size. As shown in Tab. 3, we observe consistent improvements in long video understanding compared with the video fine-tuning baseline despite an absence of any long video training data. Notably, when using the same number of 100K training samples, our hybrid data scheme improves by 6.6 points over the baseline. We hypothesize this is because the reasoning patterns within the long text can also transfer to video understanding, which exempts the need to curate specific text samples translated from vision information [27, 28].
Performance change w.r.t. frame number. We also investigate if our approach expands the context window. Larger context windows are usually beneficial since video-LLMs derived from large-context LLM backbones often accompany frame number generalization and benefit from inputting more video frames [49, 26]. However, we do not observe this trend. As shown in Fig. 8, the model fine-tuned with long synthetic samples still follows a similar pattern, that is, when performing inference beyond the video context of the training stage, the performance stays relatively stable and does not benefit from more frame input. And when the frame number exceeds the LLM context, the performance plunges to a low level. Thus, promising directions for further improvements may include continued pre-training to expand the LLM context window.
@@ 6 Conclusion
In this paper, we propose Sparrow, a data-efficient training scheme for video-LLMs, which enables training with fewer samples and achieving better video understanding performance. This method derives from our empirical findings that the low learning efficiency in data scaling may be ascribed to a limited instruction diversity in the training corpus. Thus, we design an economical data augmentation method that synthesizes video-like samples rich in instruction diversity. Comprehensive experiments demonstrate the general effectiveness and key properties of our proposed method. We hope this paper’s findings can spark more explorations of efficient training and high-quality video training corpora.
Appendix A More Implementation Details
@@ A.1 Answer Judging
A.2 Reproduction Details of Baseline Methods
Due to an inconsistent formulation between the method TOPA [27] and our proposed method, we adapt the implementation for a fair comparison. The original sample comprises a global caption for the whole video and frame-specific information. The frame-related information contains a frame-level caption and some descriptions of key objects in the frame. Thus, we design a prompt template to fit the original textual samples into the unified training format.
A real case of formatting the sample with the devised template is shown in Tab. 5.
@@ A.2 Reproduction Details of Baseline Methods
If the meaning of all options is significantly different from the answer, output Z. You should directly output a single uppercase character, such as A, B, C, D (if they are valid options) and Z, and nothing else. Here are two examples.
@@
Question: What is the main object in the image?
@@
Question: What is the main object in the image?
@@
Now here are the question, options, and the answer, you should match and give me the option letter:
@@
User:
You will be provided with some information about a video, including a global caption for the whole video, global captions for each video frame, and descriptions of key objects in each frame. Answer the questions using the information below.
Video information:
Caption: This video shows the carpentry process. At first, the person sits on a workbench and measures the length of the plank. Then, he uses a saw to cut the wooden plank into multiple pieces. Then, he uses a hammer to nail two pieces of wood together. Finally, he takes a break and drinks water, and leaves the workshop.
Frame information:
Frame 1:
Caption: the person sits on a workbench and holds a hammer in his right hand.
the person: the person is sitting on a workbench in the center of the frame. He is looking at a wooden plank in front of him.
hammer: He has a hammer in his right hand.
workbench: He is sitting on a workbench in the center of the frame.
Frame 2:
Caption: the person takes a chisel from a red toolbox and starts chiseling the plank.
chisel: There are a chisel and a hammer on the workbench in the center of the frame.
the person: the person is holding a chisel in the center of the frame.
red toolbox: There is a red toolbox in the right corner of the frame.
workbench: He is sitting on a workbench in the center of the frame.
Frame 3:
Caption: the person measures the length of the plank with a measuring tape.
the person: the person carries a pencil and a ruler in his right hand.
pencil: There is a pencil in the right hand of the person.
measuring tape: the person is holding a measuring tape in his right hand.
wooden plank: The plank is above the workbench.
Frame 8:
Caption: the person puts his hammer and chisel into the toolbox and leaves the workshop.
toolbox: the person is putting a hammer and a chisel into a toolbox in the center of the frame.
the person: the person is walking toward the door in the center of the frame.
hammer: There are a hammer and a chisel in the right hand of the person.
workshop: the person is leaving the workshop.
Question: Why does the person use a measuring tape in the third frame?
A. To hammer the plank
B. To cut the plank
C. To measure the length of the plank
D. To handle the saw
E. To fix the length of the wooden plank
Answer with the option’s letter from the given choices directly.
Assistant:
C
Table 5: An example of using the template for structuring textual samples into the training format. For simplicity, we only show part of the frame-level information.